PLEASE HELP!!!
Write an equation for the parabola with focus (4,0) and directrix y=2
I must show my work.
Thanks

Answer by scott8148(5896)

a general form for the equation of a parabola is __ (x-h)^2=4p(y-k)
__ where (h,k) is the vertex and p is the distance from the vertex to the focus

the vertex is midway between the focus and the directrix
__ in this case, (4,1)
__ so p=-1

the equation would be (x-4)^2=-4(y-1)

solving for y gives y=(-1/4)(x-4)^2 + 1